### v1

# GRAVITY: A Framework for Personalized Text Generation via Profile-Grounded Synthetic Preferences

Abstract
Personalization in LLMs often relies on costly human feedback or interaction logs, limiting scalability and neglecting deeper user attributes. To reduce the reliance on human annotations, we introduce GRAVITY (Generative Response with Aligned Values, Interests, and Traits of You), a framework for generating synthetic, profile-grounded preference data that captures users’ interests, values, beliefs, and personality traits. By integrating demographic, cultural, and psychological frameworks—including Hofstede’s cultural dimensions, Schwartz’s basic values, the World Values Survey, and Big Five OCEAN traits, GRAVITY synthesizes preference pairs to guide personalized content generation. We evaluate GRAVITY on book descriptions for 400 Amazon users, comparing it to prompt-based conditioning, standard fine-tuning, and naive synthetic pair generation. Profile-grounded synthetic data consistently improves generation, especially across multiple cultures (USA, Brazil, Japan, India), achieving over 4% higher preference gains across baselines, with user studies showing that GRAVITY outputs are preferred over 86% of the time. Our results show that scenario-grounded synthetic data can capture richer user variation, reduce reliance on costly annotation, and produce more engaging, user-centered content, offering a scalable path for LLM personalization. [^1] Code is available: https://github.com/limenlp/GRAVITY.
GRAVITY: A Framework for Personalized Text Generation via Profile-Grounded Synthetic Preferences
[…]
1 Introduction
Figure 1: Our approach GRAVITY consists of four stages: (1) User Profile Creation, (2) Custom Preference Data Selection, (3) User-Centric LLM Training, and (4) Product Description Personalization. In stage 1, we extract information about the user including explicit values such as demographic attributes: age, location, and gender and users’ interests and implicit values including their personality traits and their values and beliefs (extracted based on seed statements generated from various psychological and cultural frameworks). In stage 2, During stage 1, we generate a candidate pool of scenarios for values and beliefs based on Stage 1 using GPT-4o. We then generate a custom set of chosen/reject preference pairs for users spanning three facets: user interests, personality, and values using a combination of the Amazon Reviews dataset, personality SJTs (TRAIT and Big5Chat), and the candidate pool of generated scenarios. In stage 3, we preference tune Llama with the users’ preference pairs, and finally, in Stage 4, we generate the personalized description by prompting this tuned model with user profile attributes.
Personalization has become a critical frontier for LLMs Jang et al. (2023); Chen et al. (2024); Zhang et al. (2024b). While recent advances enable fluent and contextually relevant text generation, outputs often remain generic, overlooking individual differences in taste, style, and preference Zhang et al. (2025); Moorjani et al. (2022). In domains such as book recommendations, this gap is particularly visible—two users may value entirely different aspects of the same content, with one drawn to narrative structure and cultural context and the other to character development and personal resonance. Generic descriptions risk alienating users by failing to capture what truly matters Yunusov et al. (2024); Cai et al. (2023).
Prior work in personalization often relies on human-annotated preference data, whether through reinforcement learning from human feedback (RLHF) Kirk et al. (2024); Poddar et al. (2024), preference modeling Lee et al. (2024a); Zhong et al. (2024), or profile-conditioned prompting Zhang et al. (2024a); Lyu et al. (2023). While effective, large-scale annotation is costly and difficult to scale, limiting the breadth of user attributes that can be incorporated. Moreover, existing approaches frequently reduce personalization to narrow signals such as demographics or explicit traits, overlooking deeper dimensions of user engagement.
To combat this challenge, we introduce GRAVITY (Generative Response with Aligned Values, Interests, and Traits of You), a framework for creating synthetic, profile-grounded preference data that can be used to fine-tune LLMs for more effective personalization. As a case study, we focus on personalizing book descriptions using users from the Amazon Book Reviews dataset Hou et al. (2024). Instead of collecting explicit preference annotations, we construct synthetic preference pairs grounded in well-established psychological and cultural frameworks—including Hofstede’s cultural dimensions Hofstede (1983), Schwartz’s theory of basic values Schwartz (2012), the World Values Survey Haerpfer et al. (2024), and the Big Five OCEAN traits Goldberg (2013). This allows us to capture variation in user interests, values, beliefs, and personality traits, providing a richer basis for personalization than demographics alone. We then fine-tune Llama-3.1-8B-Instruct with Direct Preference Optimization (DPO) Rafailov et al. (2023), aligning generations with these profile-derived preferences.
We evaluate GRAVITY against several baselines, including prompt-based conditioning, standard supervised fine-tuning, and a naive DPO approach using synthetic pairs without structured profiles. Our results show that profile-grounded synthetic data consistently improves generation, achieving over 4% higher preference scores across baselines. User studies further show that GRAVITY generations are preferred over 86% of the time.
While our results do not imply that synthetic data can replace human feedback, they suggest that carefully designed, structured, scenario-based synthetic pipelines can reduce annotation needs while capturing various user attributes, offering a scalable path for aligning LLM-generated content with what users actually find engaging. Our contributions are summarized:
  • GRAVITY: We introduce GRAVITY, a multi-step framework for personalized content generation using synthetic, profile-grounded user data, capturing values, interests, and personality traits while reducing reliance on costly human annotation.
  • Profile Modeling: We develop a pipeline integrating psychological and cultural frameworks (Hofstede, Schwartz, WVS, OCEAN) to generate customized preference data that reflects diverse user profiles.
  • Comprehensive Evaluation: We assess our method using both automatic metrics and a user study, measuring preference alignment and user-perceived engagingness of the generated content.
  • Data Efficiency: We show that profile-grounded synthetic preference data yields measurable improvements over prompt-based, SFT, and naive DPO baselines.
2 Background
Personalization is central to making generated content engaging and relevant: attributes such as user interests, style preferences (especially through personality) Nguyen et al. (2018), or cultural background Matz et al. (2024) can strongly shape whether stories and characters resonate with readers Woźniak et al. (2024); Yang et al. (2023). Recent work leverages LLMs for personalization across both direct text generation and downstream applications like recommendation Jiang et al. (2023); Jang et al. (2023). Broadly, methods fall into two streams: (a) personalized generation, where models adapt outputs via prompting, user profiles, or fine-tuning Peng et al. (2024); Li et al. (2024b); and (b) LLMs for personalized tasks, such as recommendation, reasoning over user histories, or tailoring explanations Bismay et al. (2024); Shao et al. (2024). Hybrid approaches increasingly combine these, generating user-specific content that simultaneously improves alignment and task performance Shenfeld et al. (2025); Zhong et al. (2024).
Within generation, personalization extends beyond static profile conditioning to new task formulations. Work on summarization, for example, uses multi-agent pipelines to refine drafts according to user preferences Xiao et al. (2023), while headline generation has incorporated user-history modeling Song et al. (2023). Similarly, recommender systems use LLMs to expand sparse item text or construct structured user/item profiles from histories, enabling personalized reviews and explanations Lyu et al. (2023); Acharya et al. (2023).
Finally, advances in alignment and evaluation point toward more fine-grained personalization. Extensions of RLHF introduce per-user reward components or factorized preference functions Li et al. (2024b); Shenfeld et al. (2025); Poddar et al. (2024), while controllable preference vectors allow multi-objective adaptation Zhong et al. (2024). Benchmarks like LaMP Salemi et al. (2023) and newer evaluations on role-playing, user modeling, and cultural adaptation Tseng et al. (2024); Zhang et al. (2024b) reflect this growing focus on personalization as a key axis of LLM evaluation.
3 GRAVITY
In this section, we present our framework GRAVITY for generating personalized content. We first describe product description personalization (§3.1), then our user profile pipeline (§3.2) and synthetic preference data generation (§3.3). Finally, we outline user-centric LLM training (§3.4) and personalized description generation (§3.5). For evaluation, we focus on generating personalized book descriptions for 400 Amazon readers from diverse cultural backgrounds; additional details on the users and dataset are provided in §A.
3.1 Task Formulation
We formulate personalization of product descriptions as a conditional generation problem grounded in structured user profiles. Let $\mathcal{U}=\{u_{1},u_{2},\dots,u_{|\mathcal{U}|}\}$ denote the set of users, and $\mathcal{B}=\{b_{1},b_{2},\dots,b_{|\mathcal{B}|}\}$ denote the set of products. Each user $u\in\mathcal{U}$ has a historical review sequence
$R_{u}=(r_{u1},r_{u2},\dots,r_{u|R_{u}|}),$
where $r_{ui}$ is the textual review for product $b_{i}\in\mathcal{B}$.
From these reviews and basic demographic information, we curate a structured user profile $P_{u}$, which includes four components: (1) Demographics (e.g., age, gender, location), (2) Product interests, (3) Intrinsic values and core beliefs, and (4) Personality traits.
Each product $b_{n}\in\mathcal{B}$ is associated with a generic, non-personalized description $d_{n}$. Given a target user $u$ and product $b_{n}$, the goal is to generate a personalized description $\hat{B}_{n}$ that better aligns with the user’s profile $P_{u}$. Formally, this can be written:
$\hat{B}_{n}\sim P(\mathcal{D}\mid P_{u},d_{n};\theta),$
where $\mathcal{D}$ denotes the space of possible product descriptions and $\theta$ represents the model parameters. The personalized generation model can be represented as:
$\hat{B}_{n}=F_{\theta}(P_{u},d_{n}),$
where $F_{\theta}$ is the conditional generation model that incorporates both structured user information and the generic product description. The output $\hat{B}_{n}$ is a personalized description that reflects the user’s preferences and traits while preserving the content of the original product description.
Source Seed Statement Total Count
Hofstede’s Cultural Dimensions
Hofstede (1983)
Planning for the future and working toward long-term goals is more important than immediate rewards. 10
Schwartz’ Theory of Basic Values
Schwartz (2012)
It is important to me to make my own choices and decisions and think independently, even if others may disagree. 10
World Values Survey
Haerpfer et al. (2024)
I believe hard work doesn’t generally bring success—it’s more a matter of luck and connections. 130
Table 1: Example seed statements generated from various cultural and psychological constructs.
3.2 Creating User Profiles
Since personalization relies on a nuanced understanding of the user, the first step in GRAVITY is to extract salient user information and construct a structured profile. As discussed in §2, attributes such as demographics, interests, values, and personality traits can strongly influence how engaged and connected a user feels with generated content.
Each user profile in our framework consists of four key components: (1) Demographics (gender, age, location), (2) Interests, (3) Values and Beliefs, (4) Personality Traits. Demographic information is obtained directly from user-provided data when available. When gender or age is missing, we follow prior work that has shown strong performance in demographic estimation and apply trained DeBERTa models Tureyyen (2023a, b). To reduce noise and sparsity, we bin ages into three groups: young (<30), middle-aged (31–60), and senior (>60).
Interestscapture a user’s content preferences, i.e., all product categories the user might be interested in or frequently engage with. the book topics and categories they most frequently engage with. These interests can be extracted based on users’ product purchase and review history, combined with their explicitly selected preferences. Specifically, for our case study on Amazon readers, we extract users’ most frequent book genres that account for 10% or more of their reviews [^2] This threshold ensures that the selected genres reflect substantial and consistent engagement rather than incidental activity..
Values and Beliefscapture a user’s implied norms and ideals across domains such as culture, religion, ethics, politics, and society. To extract these, we adopt a multi-step LLM-based approach (see Figure 1). We first construct 150 seed value statements, drawing from established cultural theories (Hofstede’s cultural dimensions, Schwartz’s theory of basic values) and large-scale value surveys (World Values Survey). Table 1 shows representative examples. We then prompt an LLM (GPT-4o) with each seed statement and the user’s full set of product reviews, asking it to infer whether the user likely supports, opposes, or is neutral toward the statement. This yields a structured profile of the user’s values and beliefs.
Personality Traitscapture a user’s underlying preferences and tendencies, shaping both their writing style and the kinds of story elements they are likely to connect most with. To infer these traits, we use PersonalityLM Wang and Sun (2024), a fine-tuned RoBERTa classifier to predict high/low levels for each of the Big Five (OCEAN) traits. In our study, we estimate user personalities using each user’s historical book reviews.
Personalization Method Top-1 WinRate (%) Preference Gain (%) Interestingness Score Text Similarity
Original 0.75 - 3.74 -
BaseRewrite 1.25 63.5 3.81 72.8
DemoBased 3.25 68.5 3.69 72.4
UserSummary 9.25 72.0 3.98 71.0
LaMP 10.0 75.5 4.06 72.8
TriAgent 7.0 73.25 3.95 70.37
UserSFT 19.5 74.25 3.91 79.92
PrefAlign 22.25 79.0 4.02 79.48
GRAVITY (Ours) 26.75 82.5 4.03 73.2
Table 2: We compare several baselines, including existing personalization methods using prompt-based techniques and multi-agents, to our method GRAVITY. We report aggregated top-1 win rates, preference gains, interestingness scores, and text similarity scores over all selected users. Win rates, preference gains, and interestingness scores are measured with an LLM-judge (GPT-4o). We find that our method GRAVITY achieves the highest personalization metrics (↑ WinRate and Preference Gain and comparable Interestingness scores) and comparable text similarity scores to other baseline approaches. The original book description has a win rate below 1%, showing personalization methods generate more engaging descriptions.
3.3 Generating Custom Preferences
Based on users’ generated profiles, the next step in GRAVITY is to construct custom preference data that guides model adaptation toward personalized product descriptions. Rather than relying on generic preference statements, we generate scenario-based data that captures how a user’s interests, values and beliefs, and personality traits shape their judgments in concrete contexts. While the underlying pool of scenarios remains consistent across users, each profile produces a unique mapping of chosen and rejected labels, ensuring that supervision reflects their individual preferences. This scenario-driven approach provides richer, more discriminative signals than abstract descriptions, enabling the model to learn finer-grained distinctions in how different users engage with content Singh et al. (2025); Huang et al. (2023).
InterestsWe construct two types of interest-based preference data: (1) Category, which contrasts broad product categories (e.g., romance vs. biographies), and (2) Summary, which contrasts detailed book descriptions. For our study of personalized book descriptions, we begin with 382 categories from the Amazon Reviews dataset and select 3 representative, highly rated books per category. For each user, we identify their top 3–5 genres (§3.2) and, using SentenceBERT embeddings with cosine distance, select the 3 most distinct categories. We then generate Category preference pairs (chosen vs. rejected) for the user’s preferred genres. For Summary data, we form preference pairs from book descriptions of the representative titles in the user’s top and most distinct genres.
Values and BeliefsTo generate preference data about user’s values and beliefs, we first synthetically generate scenarios using GPT-4o for each seed value statement (§3.2). For each seed, GPT-4o produces 3 unique pairs of scenarios, where the first scenario strongly aligns with the statement and the second contradicts it, yielding 450 scenario pairs in total. To generate user-specific preference labels, we utilize each user profile’s extracted set of values. For a given seed statement, if the user supports it, all generated aligned scenarios are marked chosen; if the user opposes it, aligned scenarios are marked rejected [^3] Scenarios from neutral seed statements are omitted from the user’s preference dataset..
Personality TraitsWe leverage two situation-based personality questionnaires, TRAIT (scenarios) Lee et al. (2024b) and Big5Chat (dialogues) Li et al. (2024a), each containing questions with binary-level answers for the OCEAN traits. For each user, we randomly select 150 question–answer pairs from each dataset. For each pair, the answer consistent with the user’s OCEAN trait profile is labeled chosen, and the alternative labeled rejected.
3.4 User-Centric LLM Training
For our task of personalized book descriptions, we generate approximately 1,000 customized preference pairs across all three dimensions for each user (400K for user base). Using this data, we utilize Direct Preference Optimization (DPO) Rafailov et al. (2023) to generate a single personalized LLM which learns the behaviors and styles of the users in our dataset. Each training instance conditions the model on a user’s demographic, personality, values, and interests, enabling generalization across diverse profiles. At inference, we provide the target user’s profile to generate personalized book descriptions. For evaluation, we sample a highly rated title from the user’s preferred genres, excluding books they have already reviewed. Additional training details are in §4.1.
3.5 Product Description Personalization
To generate personalized descriptions, we follow prior work Santurkar et al. (2023); Hwang et al. (2023) by conditioning model responses on the user’s profile. Each profile, as described in §3.2, includes demographics, genre interests, personality traits, and personal values. Full prompt templates for description generation are provided in §C. For our task, we select a target book description to personalize from a randomly chosen highly rated title within the user’s top genres, ensuring the user has not previously purchased or reviewed the book.
4 Experiments
In this section, we outline our experimental design for personalized text generation (§4.1), including the training pipeline for user-centric LLMs, the baselines used for comparison, and evaluation strategies. We then present results across all methods (§4.2) and conduct ablations to isolate the contribution of each preference dataset to personalization performance (§4.3). Finally, we compare two training strategies for incorporating user preferences—standard supervised fine-tuning (SFT) and Direct Preference Optimization (DPO)—to assess their relative effectiveness (§4.4).
Persona Country: India
Age: Young
Gender: Female
Interests: Young Adult, Romance, Fiction
Values and Beliefs: Friends are an important part of life.; Excitement, novelty, and change are desirable in life.
Personality Traits: High Openness, High Extroversion
Original Description Our Chemical Hearts follows Henry Page, a teenager who’s never fallen in love—until Grace Town joins his class. Paired to work on the school newspaper, he discovers the excitement and challenges of first love, and the ways it can change a person.
LaMP Our Chemical Hearts follows Henry Page, a teenager who has always thought of himself as a hopeless romantic—but has never truly fallen in love. Focused on school and friendships, his life takes an unexpected turn when Grace Town joins his class. As they are paired to work on the school newspaper, Henry discovers the excitement, confusion, and intensity of first love—something any young reader navigating the ups and downs of adolescence can relate to.
PrefAlign Our Chemical Hearts tells the story of Henry Page, a teenager who considers himself a hopeless romantic but has never fallen in love. Focused on school and his future, his world changes when Grace Town joins his class. She isn’t exactly the girl he expected, but as they are paired to work together on the school newspaper, Henry begins to experience the joys and challenges of first love, discovering how unpredictable and powerful emotions can be.
GRAVITY Our Chemical Hearts follows Henry Page, a teenager who has always thought of himself as a hopeless romantic—but has never truly fallen in love. Focused on school and his friends, his life takes an unexpected turn when the intriguing and unpredictable Grace Town joins his class. As they are paired to work on the school newspaper, Henry experiences the thrill, surprises, and intensity of first love, along with the joys and challenges of navigating friendships and new experiencesperfect for any young reader who loves excitement, connection.
Table 3: Example personalized book descriptions for a simplified persona for select baselines. Our method better captures user-relevant themes compared to baselines. Highlighted texts show relevance to user’s Interests, Values and Beliefsand Personality Traits. We recommend viewing this table in full color.
4.1 Experiment Setup
Finetuning user-centric LLMs We adapt Llama-3.1-8B-Instruct Dubey et al. (2024) to user preferences using DPO with LoRA adapters Hu et al. (2022). We preference tune the model using the TRL library CarperAI (2023) with the AdamW optimizer Loshchilov and Hutter (2017), a $\beta$ value of 0.3, a learning rate of $2\times 10^{-5}$, weight decay of 0.01, a cosine learning rate schedule, and a warmup ratio of 0.05. Additional training details and prompt formats are detailed in §D.1.
Personalization strategies We benchmark against seven baselines [^4] All methods use Llama-3.1-8B-Instruct unless noted., differing in how user information is integrated: (1) Prompting: (a) BaseRewrite — rewrites the description to be more engaging (no personalization), (b) DemoBased — appends user demographics (age, gender, location), (c) UserSummary — prepends a GPT-4o–generated user profile summary, (d) LaMP Salemi et al. (2023) — retrieves and prepends the user’s most relevant past reviews, (e) TriAgent Xiao et al. (2023) — a three-stage pipeline combining BaseRewrite, GPT-4o summaries, and customized edit instructions; (2) SFT: (a) UserSFT — fine-tunes on (book description → user review) pairs with persona prompts encoding user demographics; (3) Preference data: (a) PrefAlign — uses DPO using GPT-4o-generated aligned and misaligned book descriptions, produced solely from user demographics. Details on baseline implementations are in §D.2.
Auto-evaluation metrics To assess the quality of personalized descriptions, we utilize three personalization-measuring metrics: (1) Top-1 WinRate: how often is a method’s output preferred over all other methods, (2) Preference Gain: how often is a method’s output preferred over the original description, and (3) Interestingness Score: how engaging/captivating does the user find the method’s output (based on 5-point Likert score). We evaluate personalizations using LLM-as-a-judge Zheng et al. (2023). We provide GPT-4o with a detailed user persona, generated from the user’s complete set of written reviews and and available demographics. To ensure personalized content remains faithful to the original description, we also compute cosine similarity between SentenceBERT Reimers and Gurevych (2019) embeddings of the original and generated texts. Prompt templates used for evaluation are included in §D.3.
Method Top-1 Win (%) Pref. Gain (%) Int. Score
Original 7.78 3.48
TriAgent 30.92 78.87 4.09
GRAVITY 61.30 86.73 4.15
Table 4: User Study personalization metrics (Top-1 Win Rate, Preference Gain, and Interestingness Score) based on user rankings of book descriptions: original, LaMP, and GRAVITY descriptions, aggregated over all users. Consistent with auto-evaluation metrics, we observe users prefer GRAVITY generations over baseline generations and the original description.
User study We conduct a small study with 120 participants from diverse cultural backgrounds (30 users from each of USA, Brazil, Japan, and India). For each participant, we collect demographics, top book genres, values, and personality traits. We then train a customized model for these users and generate personalized book descriptions for each participant based on popular books in their preferred genres. Participants rank three descriptions of 10 unique books (original, TriAgent, and GRAVITY) based on how engaging and interesting they find each book description. Further details on our user study design are included in §D.4.
Figure 2: Preference Gains (%) for personalized generation methods across users from four diverse countries: USA, Brazil, Japan, and India. We observe that GRAVITY yields consistently strong personalization metrics across the cultures, with gains in non-Western countries compared to current approaches (>10% increase in preference gains).
4.2 Personalization with GRAVITY
Comparisons against baselines Table 2 reports automatic evaluation metrics for all baselines and GRAVITY. Our method achieves the highest win rates (4.5% improvement) and preference gains (3.5% improvement) compared to the strongest baselines. Interestingness scores are similarly high across baselines and GRAVITY, with average ratings above 4, indicating that generated descriptions are engaging to users. Furthermore, more than 70% of the content remains semantically consistent with the original description—comparable to other baselines—demonstrating that our method preserves the original content while personalizing it.
User evaluations Table 4 show GRAVITY leads to generalizable performance gains across users from various countries and book categories. Compared to the best baseline, LaMP, we observe that users tend to prefer book descriptions generated via GRAVITY more than 61% of the time. Moreover, users also tend to prefer GRAVITY generations approximately 8% more than original book descriptions. Users also tend to find descriptions generated through GRAVITY as the most interesting.
Figure 3: Preference gains (%) across nine book genres (similar categories, clustered into fiction and non-fiction) using GRAVITY. We find that most fiction books have much higher preference gains (> 70%) as compared to non-fiction books. ($\approx$ 65%).
Cross-Cultural Personalization To evaluate robustness across diverse populations, we report preference gains for users from the USA, Brazil, Japan, and India (Figure 2). Results are averaged across each user group for the six baselines and our approach, GRAVITY. We find that GRAVITY achieves consistent improvements in all regions, with higher preferences in almost all countries. Particularly, we observe large yields (>10%) increase) in non-Western contexts such as Japan. Although Brazilians may prefer PrefAlign generations approximately 3% more than GRAVITY, we still observe comparable performances across personalization metrics (see Table 15) with GRAVITY. These results indicate that tailoring preference data to capture users’ cultural value systems and personality traits may substantially improve alignment with user’s engagement and preferences.
Feature Preference Gain (%)
All (original GRAVITY) $86.73$
- User Interests 10.34-10.34**
   - Category 6.23-6.23**
   - Summary 4.38-4.38*
- Values & Beliefs 8.92-8.92**
- Personality 8.48-8.48**
   - Openness 5.45-5.45**
   - Conscientiousness 2.35-2.35*
   - Extraversion 6.03-6.03**
   - Agreeableness 1.32-1.32
   - Neuroticism 0.98-0.98
Table 5: Ablation studies showing the impact of different user preference data (interests, values, and personality traits) on personalization performance, measured via decreases in preference gains. Subsets of Values & Beliefs data are not removed due to small size (2K–3K pairs per user). Statistical significance is assessed using paired Wilcoxon signed-rank tests: * indicates $p<0.05$, and ** indicates $p<0.01$.
Personalization Across Genres We further analyze how personalization effectiveness varies across book genres. While our method consistently improves over baselines across both fiction and nonfiction, we observe substantially higher gains for fiction titles (see Figure 3555Although the Amazon Reviews dataset has several fine-grained labels for book genres, in this analysis, we focus on a set of nine hierarchical book categories/themes.); we see preference gainst above 75% for fiction books but gains closer to 65% for non-fiction books. Our findings align with prior research which show that fiction readers tend to be more engaged with narrative elements—such as character development, emotional arcs, and imaginative settings—which naturally connect with readers’ personal values, beliefs, and personality traits Bal and Veltkamp (2013); Goyal and Mahmoud (2024). Thus, incorporating these data as training signals into user-centric models allows for better personalization. In contrast, nonfiction reading is primarily fact-oriented, emphasizing clarity, accuracy, and logical structure, which limits the potential for personalization, resulting in smaller improvements through GRAVITY.
4.3 Effects of Custom Preference Data
To understand the contribution of different components of user preference data to personalization performance, we perform an ablation study. Specifically, we remove each data subset: Interests, Values and Beliefs, and Personality Traits, and measure the resulting change in preference gains compared to the full model (Table 5) (using GPT-4o judge, see §D.3 for prompt details).
We observe that removing any preference dataset component leads to lead to statistically significant drops in preference gains, highlighting the importance of each type of user information. Notably, apart from Interests, Values and Beliefshave the largest impact, with decreases of almost 9%, while individual Personality Traitssuch as Openness and Extraversion also contribute substantially. Some traits, like Agreeableness and Neuroticism, have a smaller effect, suggesting that their influence on personalization may be more subtle.
4.4 SFT vs. DPO
While DPO enables models to directly learn user preferences and stylistic nuances, supervised finetuning (SFT) offers a simpler, more stable alternative that is less sensitive to hyperparameter choices and requires less training resources. To study this trade-off, we evaluate a setting where models are taught about a user’s preferred book genres, values, and personality traits through SFT. For each user, we construct binary training examples of the form: given demographic attributes and a scenario, the user would most likely $A$, where $A$ corresponds to the chosen response from each preference pair.
As in our original approach, we train Llama-3.1-8B-Instruct, using LoRA for 5 epochs using a learning rate of $10^{-4}$ with GPT-4o as our evaluation judge. Table 6 reports results across baselines and GRAVITY: SFT and DPO.
Method Top-1 Win (%) Pref. Gain (%) Int. Score
BaseRewrite 1.0 63.5 3.81
DemoBased 3.0 68.5 3.69
UserSummary 4.5 72.0 3.98
LaMP 4.75 75.5 4.16
TriAgent 4.25 73.25 3.95
UserSFT 17.0 74.25 4.04
PrefAlign 20.75 79.0 4.24
GRAVITY (SFT) 20.25 72.58 3.98
GRAVITY (DPO) 24.5 82.67 4.10
Table 6: Personalization metrics with SFT on GRAVITY data. SFT improves engagement (comparable to PrefAlign and UserSFT method results), while DPO achieves the strongest gains in win rate, preference, and interestingness.
Overall, our findings suggest that incorporating user-specific data (interests, values, and personality) enhances personalization across methods. However, DPO more effectively captures fine-grained preferences, leading to the highest improvements in alignment and engagement for win rate, preference gain, and user interestingness.
5 Conclusion
In this work, we presented GRAVITY, a framework for generating personalized text by leveraging synthetic, profile-grounded preference data derived from demographics, interests, values, beliefs, and personality traits. By integrating psychological and cultural frameworks, our approach enables LLMs to align content with diverse user profiles and capture meaningful variation in individual preferences.
Applied to personalized Amazon book descriptions, GRAVITY demonstrates that scenario-driven preference data can effectively guide model behavior, with deeper attributes such as values and personality providing measurable gains beyond demographics and interests. These results highlight the potential of structured synthetic preference data as a scalable and interpretable approach to user-centric LLM personalization.
6 Limitations
In this work, we approximate user attributes in order to build a synthetic preference dataset. While this approach reduces the need for costly annotation, it may not guarantee exact alignment with an individual’s intrinsic value systems, cultural beliefs, or personality traits. Prior work has shown that such attributes can be estimated through model prompting and training, but we acknowledge that profile inference introduces uncertainty and potential noise. To partly mitigate this, we conducted a user study (see §D.4) to validate whether personalized generations align with the participants’ self-reported preferences and profile attributes.
Another limitation lies in cross-cultural generalization. Although we ground GRAVITY in established cultural psychology frameworks such as Hofstede’s dimensions, Schwartz’s values, and the World Values Survey, these instruments may not fully capture the richness and evolving nature of cultural identity. Our results show meaningful gains across different cultural groups, but personalization may still miss subtle or context-dependent user preferences that fall outside of these frameworks.
Finally, this work focuses on book description personalization using Amazon reviews as a controlled testbed for exploring the effectiveness of GRAVITY. The main goal of this work is to demonstrate the feasibility of using synthetic, profile-grounded preference data for personalization, but future work can test its applicability across broader datasets and domains, such as news, healthcare, and education, where personalization challenges may differ.
Acknowledgments
We would like to thank researchers from USC LIME and HUMANS labs for their continuous feedback in the formulation and setting of our work. This research was supported in part by the NSF, under Award Number 2331722. Priyanka Dey was also supported by the 2025 USC Capital One Fellowship.
Appendix A Dataset
As a case study for evaluating GRAVITY, we utilize the Amazon Reviews Dataset collected in 2023 and contains over 571.54M reviews spanning May. 1996 to Sep. 2023. This dataset consists of product reviews from a wide range of categories, including books, clothing, and electronics. For our study, we focus on users who are avid readers and presenting them with personalized book descriptions. The original dataset consists of approximately 22.5M total book reviews from over 5M users. In our study, we select a small set of users (400) from 4 varying countries: United States, Brazil, India, and Japan and those who have rated at least 50 books. In Table 7, we present more details on the randomly selected users.
Country Total Users Top Book Genres Avg. # of Reviews
USA 100 Young Adult, Romance, History, Business 79
Brazil 100 Historical Fiction, Philosophy, Art History, Motivational 61
India 100 Thriller/Mystery, Historical Fiction, Biographies/Memoirs, Self-Improvement 56
Japan 100 Poetry, Folk literature, Engineering, Biographies/Memoirs 53
Table 7: Details of users randomly selected for our case study of book description personalization (auto-evaluation).
Appendix B Custom Preference Generation
In this section, we provide additional details for custom preference generation for each user.
InterestsWe construct user interest data from each user’s top book genres. Across 400 users, we identify 382 unique categories (e.g., Mythology, Classical Literature, Self-Improvement). To identify the most distinct categories, we embed all category names using all-MiniLM-L6-v2 (a Sentence-BERT model) and compute pairwise cosine similarities. For each user, we compare their top category with all others and select the three least similar (i.e., most dissimilar) categories. We then form Category pairs by combining each user’s top genres with these inferred distinct categories. To construct Summary pairs, we retrieve three highly rated books from both the user’s top and distinct categories, and pair them using their original book descriptions. This process yields roughly 90–240 Interests pairs per user.
Values and BeliefsTo construct values preference data, we focus on generating a set of 150 seed statements from various cultural frameworks on a wide variety of topics including culture, society, ethics, religion, politics, and morals. Table 8 consists of example statements generated from the various frameworks and topics.
Source Seed Statement Topic
Hofstede’s Cultural Dimensions
Hofstede (1983)
Competition and achievement are more valued than cooperation and care. Culture
It is better to rely on clear rules and traditions than to face unpredictable situations without guidance. Ethics
A person’s identity comes from belonging to their group or community, not from individual achievements alone. Society
Authority should be respected, and decisions from leaders are not to be questioned Politics
Planning and persistence today are essential to secure the success of future generations. Morals
Schwartz’ Theory of Basic Values
Schwartz (2012)
Life is meant to be full of adventure and new experiences. Culture
People should be free to think, choose, and act for themselves. Ethics
Success should be measured by what a person achieves through their abilities. Society
A good society protects its people through order, stability, and safety Politics
Respecting traditions and the beliefs of past generations is a moral duty. Morals
World Values Survey
Haerpfer et al. (2024)
Leisure time is important in life. Culture
I trust people from different nationalities than my own. Ethics
If a woman earns more money than her husband, it’s almost certain to cause problems. Society
On the whole, men make better political leaders than women do. Politics
It is a duty towards society to have children. Morals
Whenever science and religion conflict, religion is always right Religion
Table 8: Additional Values and Beliefs seed statements derived from various cultural frameworks including Hofstede’s Cultural Dimensions, Schwartz’ Theory of Basic Values, and World Values Survey.
After curation of these seeds, we generate pairs of scenarios to curate a set of value/beliefs scenario candidate pool. We use the following prompt template to generate 3 unique scenarios for each seed: For the following statement: {seed statement}, please generate 3 pairs of role-playing scenarios, where the first scenario illustrates the statement and the second contradicts it. Please limit each scenario to 2-3 sentences and do not repeat scenarios. Table 9 presents example generated scenario pairs, generated from GPT-4o.
Topic Seed Statement Example Scenario Pair
Culture Leisure time is important in life. Support: After finishing work, Maria switches off her laptop and joins her friends for a long walk in the park. She tells them she believes relaxation and fun are just as essential as hard work.
No Support: John works late every evening, skipping outings and hobbies to finish more projects. When asked about taking a break, he insists leisure is a waste of time compared to productivity.
Ethics I trust people from different nationalities than my own. Support: At an international conference, Aisha gladly shares her research data with a colleague from another country, confident they will use it responsibly. She says collaboration works best when people trust each other across borders.
No Support: During a group project, Mark refuses to let a teammate from abroad handle key parts of the work. He mutters that people from other countries can’t be relied on the same way as his own nationals.
Society If a woman earns more money than her husband, it’s almost certain to cause problems. Support: At an international conference, Aisha gladly shares her research data with a colleague from another country, confident they will use it responsibly. She says collaboration works best when people trust each other across borders.
No Support: During a group project, Mark refuses to let a teammate from abroad handle key parts of the work. He mutters that people from other countries can’t be relied on the same way as his own nationals.
Politics On the whole, men make better political leaders than women do. Support: During a classroom debate, Alex argues that history shows most successful leaders have been men, so they are naturally better suited for politics. His classmates nod in agreement, saying women are better at supporting roles.
No Support: At a community meeting, Priya points to examples of female leaders who successfully managed crises with empathy and strength. She argues that leadership depends on skill and vision, not gender, and the audience applauds.
Morals It is a duty towards society to have children. Support: At a family gathering, Amina explains to her cousin that she and her husband are eager to start a family soon. She says raising children is their responsibility to continue the community and care for the next generation.
No Support: During a conversation with friends, Daniel mentions he and his partner decided not to have children. He argues that contributing to society can also mean supporting others, mentoring youth, or focusing on community work.
Religion Whenever science and religion conflict, religion is always right. Support: During a family debate about evolution, Fatima says she accepts the religious creation story over scientific explanations. She explains that when science and faith disagree, faith must guide the truth.
No Support: In biology class, Alex learns about the Big Bang and accepts the evidence despite his church teaching otherwise. He tells his classmates that scientific proof carries more weight than religious doctrine in such conflicts.
Table 9: Example scenario pairs for different seed statements across topics. Each seed is expanded into 3 pairs of scenarios, where the first aligns with the statement and the second contradicts it.
To determine each user’s values and beliefs list, we prompt the model with the following prompt:
You will be provided with a set of reviews a user has written as well as demographic attributes including age, gender, and country. For the following statement: {seed statement}, please select ONE of {’support’, ’no support’, ’neutral’} based on how well the statement reflects the user’s beliefs. Here are the list of reviews: {reviews} and demographic information: {demographics}.
After generation of each user’s value and belief system, preference labels are auto-mapped. Each user has up to 450 pairs of values/belief preference pairs.
Personality TraitsTo generate personality preference pairs, we utilize two existing datasets: TRAIT Lee et al. (2024b) and Big5Chat Li et al. (2024a). In Table 10, we summarize main statistics from these two datasets. We randomly sample 300 total questions (60 questions for each OCEAN trait, 30 from each dataset) [^6] As TRAIT contains 2 answers for each binary label answer, we randomly select one of the answers from each level to generate a preference pair Table 11 contains example questions from each dataset.
Dataset Total Qs (Big5) Scenario Format Answer Format
TRAIT Lee et al. (2024b) 5,000 (400 per trait) Situation-based Binary (2 answers per low/high)
Big5Chat Li et al. (2024a) 100,000 (20K per trait) Dialogue-based Binary
Table 10: Dataset statistics for LLM-based psychometric SJTs.
Dataset Trait Scenario Pair
TRAIT Openness Q: I go to the country fair, which is renowned for its vibrant display of local arts and crafts, including paintings, sculptures, and handmade jewelry. The fair is bustling with people of all ages, and the air is filled with the scent of fresh food and the sound of live music. How should I spend my time at the country fair to make the most of this experience?
High: Explore each artist’s booth to discover unique pieces and engage with the creators about their inspirations.
Low: Visit the top-rated food stalls first to enjoy the local flavors that everyone recommends.
Big5Chat Conscientiousness Q: Hey Noelani, I just got the most insane idea for a new stunt! We should totally try it out and see how crazy we can get! I know you’ll love it, and I can already imagine you yelling at the top of your lungs!
High: Whoa, hold up, let’s not get ahead of ourselves. I’m all for trying new things, but we need to assess the risks and make sure we’re not putting ourselves or others in harm’s way. Let’s break it down, analyze the potential consequences, and then decide if it’s worth pursuing.
Low: WOOHOO, YEAH, LET’S DO IT! I don’t care if it’s crazy or not, it sounds like a blast! I’m so down, let’s just go for it and see what happens! Who cares about the consequences, we’ll figure it out later!
TRAIT Extraversion Q: I am scheduled to lead a team meeting at work where we need to brainstorm and decide on our next project’s direction. How should I approach this meeting to make it productive and inclusive?
High: Encourage open dialogue by asking each team member to pitch at least one idea, ensuring everyone contributes and feels involved.
Low: Prepare a structured agenda with clear objectives and distribute it before the meeting to allow team members to prepare their thoughts.
Big5Chat Agreeableness Q: Hey, Gillian! Need a hand with something? You looked a bit lost when you got out of your car. Is everything okay?
High: Thanks so much, Eathan! I appreciate your help. I was just trying to get my bearings. I’m a bit turned around. Could you point me in the direction of the nearest coffee shop? I need a caffeine boost before my meeting. Low: What’s it to you, Eathan? I don’t need your help. I can handle my own problems. And what’s with the small talk? You’re not getting anything out of this, so just leave me alone.
TRAIT Neuroticism Q: I have been invited to perform at a local community event, but as the day approaches, I become increasingly nervous about performing in front of a crowd. How should I manage my anxiety and prepare for the performance?
High: You might consider preparing some back-up plans or escape strategies in case the performance feels too overwhelming when you’re on stage.
Low: Focus on the excitement of performing and the joy it brings, rather than the nerves; see this as a chance to grow and have fun.
Table 11: Example scenario pairs for personality preference pairs based on LLM-based psychometric SJTs.
Appendix C Product Description Personalization
To generate the final personalized book description for each user, we use a templatized prompt based on their demographics:
You are a {age} {gender} from {country} with {interests}. You have {personality traits} and {values}. Please generate a personalized description of {book} with the original description: {description}.
Appendix D Experiment Setup
D.1 Finetuning User-Centric LLMs
To finetune Llama-3.2-3B-Instruct, we use DPO and 4-bit quantized LoRA. We preference-tune a single model for all users using the TRL library. We use a final learning rate of: $2\times 10^{-5}$ and $\beta$ value of: 0.3. We use a batch size of 64 with a gradient accumulation step size of 2 and 3 epochs with early stopping criterion. We train our model on 2 A6000s with DeepSpeed Rajbhandari et al. (2020) which takes approximately 11 hours. Table D.2 contains our training prompts for preference aligning the model.
D.2 Personalized Prompting Strategies
We employ three kinds of baselines: (1) Prompting-based, (2) SFT, and (3) Preference-Based. For prompting based models, we allow the model to generate 6-8 sentences to keep length similar to original descriptions. Table 12 provides details on each approach as well as prompt templates.
For the LaMP-based approach, we first identify the most relevant book descriptions to the target book recommendation. To do this, we first encode all the descriptions of the books a user has reviewed into dense vector representations using a SentenceBERT (all-MiniLM-L6-v2) model. Each description is thus mapped to a high-dimensional embedding capturing its semantic content. Given a target book, we similarly encode its description into the same embedding space and retrieve the top-5 most similar books based on cosine similarity. For efficient similarity search, we leverage FAISS Douze et al. (2024) to index and query the embeddings. Based on the extracted similar books, we prepend the user’s reviews for these books as context in the generation prompt.
For our SFT approach, we finetune Llama using 4-bit quantized LoRA for 5 epochs with early stopping criterion. We train with a learning rate of $2\times 10^{-4}$. For our naive preference-based method, we use DPO to align the model to preferences. For each user, we generate a total of $N$ preference pairs (aligned and misaligned book descriptions) based on the books they have reviewed; $N$ refers to the total number of books the user has reviewed. For users with less than 100 reviewed books, we sample 100N100-N books from their top genres and also generate preference pairs for these books. Thus, each user has at a minimum 100 preference pairs. We train the model for a total of 3 epochs with early stopping. We use a learning rate of $2\times 10^{-5}$. For both training baselines: SFT and DPO methods, we utilize the TRL library, AdamW optimizer, weight decay of 0.01, a cosine learning rate scheduler, and a warmup ratio of 0.05. We also utilize DeepSpeed and 2 A6000’s to train the models.
Baseline Prompt Template/Training Example
BaseRewrite Personalization Prompt: Please generate a more engaging and interesting description for this book: book with this description: {description}.
DemoBased Personalization Prompt: You are a {age} {gender} from {country}. Please generate a more engaging and interesting description for this book: {book} with this description: {description}.
UserSummary User-Summary Generation Prompt: Please generate a summary of the user based on their historical reviews: $r_{1}$, $r_{2}$, …, $r_{n}$]. Personalization Prompt: Based on this user summary (user_summary), please generate a more engaging and interesting description for this book (book) based on the following description: description.
LaMP Personalization Prompt: Based on these user reviews: $r_{1},r_{2},r_{3},r_{4},r_{5}$], please generate a more engaging and interesting description for this book (book) based on the following description: description.
TriAgent First Generation Prompt: Please generate a more engaging and interesting description for this book: {book} with this description: {description}.
Edit Instructions Prompt: Based on this user summary {user_summary} and this personalized book description: {personalized_description}, please generate a set of suggested edits to make the description more engaging and interesting for the user.
Final Generation Prompt: Based on this user summary {user_summary} and these suggested edits, please generate a more engaging and interesting description for this book: {book} with this description: {personalized_description}.
UserSFT Training Prompt: You are a {age} {gender} from {country}. You recently read the book: {book} with the following description: {description} and this was your review:
Training Output: {review}
PrefAlign Aligned Description Generation Prompt: You are a {age} {gender} from {country}. Please generate a more engaging and interesting description for this book: {book} with this description: {description}.
Misaligned Description Generation Prompt: You are a {age} {gender} from {country}. Please generate a less engaging or interesting description for this book: {book} with this description: {description}.
DPO Training Prompt: You are a {age} {gender} from {country}. Which book description is more engaging and interesting?
Chosen: {Aligned Generated Description}
Rejected: {Misaligned Generated Description}
Final Description Generation Prompt: You are a {age} {gender} from {country}. Please generate a more engaging and interesting description for this book: book with this description: {description}.
GRAVITY User Value Survey: Given the user’s set of values and beliefs, please construct a short paragraph (max. 6 sentences), summarizing the user’s values on culture, ethics, society, politics, politics, morals, and religion.
DPO Training Prompt: You are a {age} {gender} from {country}. You have the following values: {user_value_summary} and personality traits: {traits}. {preference_prompt}
Chosen: {preference_prompt_chosen}
Rejected: {preference_prompt_rejected}
Final Description Generation Prompt: You are a {age} {gender} from {country}. You have the following values: {user_value_summary} and personality traits: {traits}. Please generate a more engaging and interesting description for this book: book with this description: {description}.
Table 12: All baselines with detailed prompt templates and training examples.
D.3 Auto-Evaluation Metrics
For evaluation, we utilize GPT-4o as an LLM-judge. To simulate a user, we generate a persona prompt that includes their demographics (age, gender, and country) as well as a user summary, which is generated by GPT-4o from all of the user’s historical reviews. For each user, we generate a total of five personalized book descriptions, resulting in 2,000 generations per method. The LLM-judge is then provided with all personalized generations and asked to rank them according to engagement and interest using the following prompt:
You are a {age} {gender} from {country}. {user_summary}. Please carefully read each of the following book descriptions and provide a ranking of how engaging and interesting you find each description: [description_0777This refers to the original book description, description_1, description_2, description_3, description_4, description_5, description_6, description_7].
We then calculate the WinRate for each method by computing the percentage of times the model achieves first place compared to all other methods across all users and generations. We also compute Preference Gain by calculating the percentage of times the model’s output is preferred over the original description. For measuring Interestingness, we prompt the model one description at a time using the following Likert scale (1-5): 1 = not very interesting/engaging, 5 = very interesting/engaging.
You are a {age} {gender} from {country}. {user_summary}. Please carefully read the following book description and rate how engaging and interesting you find the description from 1 to 5: {description}.
D.4 User Study Design
We conduct a multi-part user study on Prolific to evaluate personalized generation across diverse cultural backgrounds. We recruit 30 participants each from the United States, Brazil, Japan, and India. In the first phase (see Figure 4), we extract user attributes including gender and age group (as defined in §3.2). To capture users’ literary preferences, participants select their top three book genres from a list of nine categories [^8] We adopt the hierarchical genre structure shown in Figure 3..
Figure 4: Example questions from first stage in user study: data collection. Along with standard demographics, we extract user interests i.e. ranking top 3 genres, value systems, and personality.
We further measure users’ personality traits using the Mini-IPIP questionnaire Donnellan et al. (2006), a validated 20-item Likert-scale instrument for assessing the Big Five (OCEAN) traits. To estimate users’ values and beliefs, we adapt our 150-item seed statement bank, selecting a concise subset of 10 statements, across various categories: culture, ethics, society, politics, morals, and religion, to minimize participant fatigue. We then prompt GPT-4o to infer users’ likely responses to the remaining statements based on their annotated subset:
Given a user’s value system: {annotated seed statements}, would the user agree, disagree, or be neutral to this statement: {additional seed statement}?
To validate GPT-4o’s value inference accuracy, in Step 2 of our study (Figure 5), each participant is presented with five additional seed statements and asked to indicate if they agree with the GPT-4o annotation label. Across all four nations, GPT-4o achieves an average accuracy of 84%, indicating that the model can effectively infer user value systems from limited supervision (10 examples per user).
Figure 5: Example questions from second stage in user study: user values verification. Users are given additional values seed statements and asked to verify whether they agree with GPT-4o annotations.
In the final phase of our study, each participant is shown ten triplets of book descriptions: the original, one generated by GRAVITY, and one by a strong baseline model (TriAgent), which does not rely on additional user or synthetic data. Participants are asked to (1) rank the descriptions by engagement and interest, (2) rate each on a 1–5 Likert scale for interestingness, and (3) note portions they find particularly appealing (Figure 6). Table 13 illustrates an example annotation. Overall, users consistently report that GRAVITY outputs align more closely with their cultural norms, personal experiences, and value systems, underscoring the importance of integrating culture and values into personalization.
Figure 6: Example questions from second stage in user study: user values verification. Users are given additional values seed statements and asked to verify whether they agree with GPT-4o annotations.
User Details User Highlights and Rationales
Country: Brazil Age: Middle-Aged Gender: Female Value: strong supporter of women’s rights and abortion Highlight: When Nnu discovers she must make a life-changing choice, she confronts the pressure of family expectations and societal judgment, ultimately finding the courage to take control of her own body and destiny, standing firm in the belief that her choices belong to her alone. Rationale: Feminism is something that I have been passionate about for many years now. I felt like it captured exactly what it’s like to make your own choices, even when everyone around you has opinions.
Country: USA Age: Senior Gender: Female Value: life should be a mixture of fun and work Highlight: She discovered that the secret to happiness wasn’t escaping responsibility, but weaving laughter and discovery into every day’s obligations. Rationale: I think the same way as the main character here, so it probably resonated most with me.
Country: India Age: Middle-Aged Gender: Male Personality: high Neuroticism and low Extraversion Highlight: Alone in his dimly lit apartment, Ronald felt the walls closing in as the thought of his feelings became louder and he began pondering how he would escape from… Rationale: The way this was written somehow resonated well with me, maybe because I think I might respond in a similar manner in a given situation.
Country: Japan Age: Senior Gender: Male Value: supporter of culture and religion Personality: high Openness Highlight: Learn more about various Aztec festivals, including Toxcatl, one of the largest festivals devoted to Tezcatlipoca, the god of the… Rationale: This was particularly interesting to me which none of the other descriptions highlighted that well because I’m very interested in different cultures and this one highlighted an intriguing old celebration.
Table 13: Example highlights users selected from GRAVITY generations and reasons/rationales for why the description was more engaging and interesting to them. Many users highlight notions from their cultural, intrinsic values/beliefs, and personality/personal experiences.
Appendix E Additional Results
In this section, we present personalization metrics (Top-1 WinRates, Preference Gains, Interestingness Scores) split by the Amazon readers’ country. We find that split across various countries, GRAVITY generations can yield substantial yields especially in Non-Western settings (see Tables 14, 15, 17, and 16).
Personalization Method Top-1 WinRate (%) Preference Gain (%) Interestingness Score
Original 0.75 - 3.65
BaseRewrite 1.75 68.0 3.72
DemoBased 3.5 76.25 3.60
UserSummary 9.75 74.0 3.92
LaMP 13.25 79.75 4.00
TriAgent 7.5 78.25 3.88
UserSFT 21.25 77.75 3.85
PrefAlign 21.0 79.5 3.95
GRAVITY (Ours) 26.25 80.75 3.97
Table 14: Automatic personalization metrics (GPT-4o evaluated) for American Amazon readers.
Personalization Method Top-1 WinRate (%) Preference Gain (%) Interestingness Score
Original 0.5 - 3.75
BaseRewrite 1.0 68.0 3.80
DemoBased 3.75 68.5 3.72
UserSummary 10.25 73.0 3.99
LaMP 5.5 78.75 4.05
TriAgent 7.75 76.25 3.94
UserSFT 20.5 75.5 3.90
PrefAlign 24.5 81.25 4.02
GRAVITY (Ours) 27.25 78.5 4.03
Table 15: Automatic personalization metrics (GPT-4o evaluated) for Brazil Amazon readers
Personalization Method Top-1 WinRate (%) Preference Gain (%) Interestingness Score
Original 0.75 - 3.80
BaseRewrite 0.75 60.25 3.85
DemoBased 4 70.75 3.72
UserSummary 7.0 68.75 4.03
LaMP 12.5 70.5 4.08
TriAgent 6.5 69.25 3.97
UserSFT 17.75 71.25 3.92
PrefAlign 21.25 74.5 4.04
GRAVITY (Ours) 28.0 86.25 4.10
Table 16: Automatic personalization metrics (GPT-4o evaluated) for Japanese Amazon readers.
Personalization Method Top-1 WinRate (%) Preference Gain (%) Interestingness Score
Original 1.0 - 3.76
BaseRewrite 1.5 60.75 3.87
DemoBased 1.75 59.25 3.74
UserSummary 10.0 72.5 4.06
LaMP 8.75 70.25 4.11
TriAgent 6.25 72.25 3.96
UserSFT 18.5 74.5 3.97
PrefAlign 19.5 74.25 4.08
GRAVITY (Ours) 28.25 83.75 4.10
Table 17: Automatic personalization metrics (GPT-4o evaluated) for Indian Amazon readers.